OpenStudy (anonymous):

The line of symmetry for the quadratic equation y = ax 2 - 8x - 3 is x = 2. What is the value of "a"? a.)-2 b.)-1 c.)2

OpenStudy (anonymous):

\[y=a \left( x^2-\frac{ 8x }{ a }+\left( \frac{ 4 }{ a } \right)^2 \right)-\frac{ 16 }{ a }-3\] \[=a \left( x-\frac{ 4 }{ a } \right)^2-\frac{ 16 }{ a }-3\] \[y+\frac{ 16 }{ a }+3=\left( x-\frac{ 4 }{ a } \right)^2\] axis of symmetry is \[x-\frac{ 4 }{ a }=0~or~ x=\frac{ 4 }{ a }\] but x=2~is axis of symmetry Hence \[2=\frac{ 4 }{ a }\] a=?